# Learning under Imitative Strategic Behavior with Unforeseeable Outcomes

### v1

@@ 4.2 Preferences as (dis)incentives for manipulation
Compared to the original strategic policy $\theta^{*}$, decision-makers by adjusting preferences can disincentivize manipulation (i.e., $P_{M}(\theta)$ decreases) under certain scenarios. Specifically,
@@
    When either of the followings is satisfied, and the decision-maker adjusts preferences by increasing $k_{1}$:
@@ 4.3 Preferences shape algorithmic fairness
The threshold shifts under adjusted preferences further allow us to compare these policies against a certain fairness measure. In this section, we consider strategic individuals from two social groups $\mathcal{G}_{a},\mathcal{G}_{b}$ distinguished by some protected attribute $S\in\{a,b\}$ (e.g., race, gender). Similar to Zhang et al. (2019, 2020b, 2022), we assume the protected attributes are observable and the decision-maker uses group-dependent threshold policy $\pi_{s}(x)=\textbf{1}(x\geq\theta_{s})$ to make decisions about $\mathcal{G}_{s},s\in\{a,b\}$. The optimal threshold for each group can be found by maximizing the utility associated with that group: $\max_{\theta_{s}}\mathbb{E}[R(D,Y)|S=s]$.
Fairness measure. We consider a class of group fairness notions that can be represented in the following form (Zhang et al., 2020a; Zhang and Liu, 2021):
@@
Thm. 4.6 identifies all scenarios under which a decision-maker can simultaneously promote fairness and disincentivize manipulation by simply adjusting $k_{1},k_{2}$. Otherwise, it is not guaranteed that both objectives can be achieved at the same time, as stated in Corollary 4.7. Importantly, Thm 4.6 sheds light on how the decision-maker can make explainable and socially responsible decisions under the unforeseeable strategic individual behavior: instead of adding separate regularizers to prevent manipulation or promote fairness, we show that the decision-maker may simply adjust their preferences in an interpretable way to incentivize improvement and promote fairness at the same time.
@@
The results above assume the decision-maker knows $q,\epsilon$ precisely. In practice, these parameters may need to be estimated empirically. In App. B.5, we further provide an estimation procedure for the parameters, enabling the decision-maker to design a policy starting with the population data.
@@ 5 Experiments
We conduct experiments on both synthetic Gaussian data and FICO score data (Hardt et al., 2016b).
Figure 3: $P_{M}(\theta)$ of Caucasian and African American.
FICO data (Hardt et al., 2016b). FICO scores are widely used in the US to predict people’s credit worthiness. We use the preprocessed dataset containing the CDF of scores $F_{X|S}(x|s)$, qualification likelihoods $P_{Y|XS}(1|x,s)$, and qualification rates $\alpha_{s}$ for four racial groups (Caucasian, African American, Hispanic, Asian). All scores are normalized to $[0,1]$. Similar to Zhang et al. (2022), we use these to estimate the conditional feature distributions $P_{X|YS}(x|y,s)$ using beta distribution ${Beta}(a_{ys},b_{ys})$. The results are shown in Fig. 9. We assume the improved feature distribution $P^{I}(x)\sim{Beta}\big{(}\frac{a_{1s}+a_{0s}}{2},\frac{b_{1s}+b_{0s}}{2}\big{)}$ and $C_{M}-C_{I}\sim\mathcal{N}(0,0.25)$ for all groups, under which Assumption 2.2 and 2.1 are satisfied (see Fig. 8). We also considered other feature/cost distributions and observed similar results. Note that for each group $s$, the decision-maker finds its own optimal threshold ($\theta^{*}_{s}$ or $\theta^{*}_{s}(k_{i})$ or $\widehat{\theta}^{*}_{s}$) by maximizing the utility associated with that group, i.e., $\max_{\theta_{s}}\mathbb{E}[R(D,Y)|S=s]$.
@@
Figure 4: Impact of adjusted preferences (FICO data)
We also examine the individual best responses. Fig. 3 shows the manipulation probability $P_{M}(\theta)$ as a function of threshold $\theta$ for Caucasians (blue) and African Americans (orange) when $q=0.3,\epsilon=0.5$. For both groups, there exists a unique $\theta_{max}$ that maximizes the manipulation probability. These are consistent with Thm. 2.3. We also indicate the manipulation probabilities under original strategic optimal thresholds $\theta_{s}^{*}$; it shows that African American has a higher manipulation probability than Caucasians. Similar results for Asian and Hispanic are shown in Fig. 12.
Note that the scenario considered in Fig. 3 satisfies the condition 1.(ii) in Thm. 4.5, because the original strategic $\theta_{s}^{*}<\theta_{max}$ for both groups. We further conduct experiments in this setting to evaluate the impacts of adjusted preferences. We first adopt EqOpt as the fairness metric, under which $\mathbb{E}_{X\sim P_{s}^{\mathcal{C}}}[\textbf{1}(X\geq\widehat{\theta})]=F_{X|YS}(\theta|1,s)$ and the unfairness measure of group $\mathcal{G}_{a},\mathcal{G}_{b}$ can be reduced to $\big{|}F_{X|YS}(\theta|1,a)-F_{X|YS}(\theta|1,b)\big{|}$. Experiments for other fairness metrics are in App. C.1. The results are shown in Fig. 4, where dashed red and dashed blue curves are manipulation probabilities under non-strategic $\widehat{\theta}^{*}$ and strategic $\theta^{*}(k_{1})$, respectively. Solid red and solid blue curves are the actual utilities $U(\widehat{\theta}^{*})$ and $U(\theta^{*}(k_{1}))$ received by the decision-maker. The difference between two dotted green curves measures the unfairness between Caucasians and African Americans. All weights are normalized such that $k_{1}=1$ corresponds to the original strategic policy, and $k_{1}>1$ indicates the policies with adjusted preferences. Results show that when condition 1(ii) in Thm. 4.5 is satisfied, increasing $k_{1}$ can simultaneously disincentivize manipulation ($P_{M}$ decreases with $k_{1}$) and improve fairness. These validate Thm. 4.5 and 4.6.
Table 2: Comparison between three types of optimal thresholds (FICO data). For values of “Utility" and “$P_{M}$", the left value in parenthesis is associated with Caucasian, while the right is for African American.
@@
Table 2 compares the non-strategic $\widehat{\theta}^{*}$, original strategic ${\theta}^{*}$, and adjusted strategic ${\theta}^{*}(k_{1})$ when $k_{1,c}=k_{1,aa}=1.5$. It shows that decision-makers by adjusting preferences can significantly mitigate unfairness and disincentivize manipulation, with only slight decreases in utilities. Results for Asians and Hispanics are in Table 5. We also present more results when these parameters are noisy in App. B.5.
@@
Gaussian Data. We also validate our theorems on synthetic data with Gaussian distributed $P_{X|YS}$ in App. C.2. Specifically, we examined the impacts of adjusting preferences on decision policies, individual’s best response, and algorithmic fairness. As shown in Fig. 21, 22 and Table 688, these results are consistent with theorems, i.e., adjusting preferences can effectively disincentivize manipulation and improve fairness. Notably, we considered all three scenarios in Thm. 4.5 when condition 1.(i) or 1.(ii) or 2 is satisfied. For each scenario, we illustrate the individual’s best response $P_{M}$ in Fig. 21 and show that manipulation can be disincentivized by adjusting preferences, i.e., increasing $k_{1}$ under condition 1.(i) or 1.(ii), or increasing $k_{2}$ under condition 2.
6 Societal Impacts & Limitations
@@
This paper proposes a novel probabilistic framework and formulates a Stackelburg game to tackle imitative strategic behavior with unforeseeable outcomes. Moreover, the paper provides an interpretable decomposition for the decision-maker to incentivize improvement and promote fairness simultaneously. The theoretical results depend on some (mild) assumptions and are subject to change when $\epsilon,q,C_{M},C_{I}$ change. Although we provide a practical estimation procedure to estimate the model parameters, it still remains a challenge to estimate model parameters accurately due to the expensive nature of doing controlled experiments. This may bring uncertainties in applying our framework accurately in real applications.
Impact Statement
@@ A.1 Strategic classification
Generally, strategic behaviors can cause feature and label distribution of individuals to shift, which have long been closely related to concept drift (Lu et al., 2018), preference shift (Carroll et al., 2022), and algorithm recourse (Karimi et al., 2022). Strategic classification has been extensively studied since (Hardt et al., 2016a) formally modeled the interaction between individuals and a decision maker as a Stackelberg Game, and proposed a framework for strategic classification. While taking the individuals’ best response into account, the decision maker can make the optimal decision by anticipating strategic manipulation. During recent years, more complex models on strategic classification have been proposed (Ben-Porat and Tennenholtz, 2017; Dong et al., 2018; Braverman and Garg, 2020; Jagadeesan et al., 2021; Izzo et al., 2021; Ahmadi et al., 2021; Tang et al., 2021; Zhang et al., 2020b, 2022; Eilat et al., 2022; Liu et al., 2022; Lechner and Urner, 2022; Chen et al., 2020b). Ben-Porat and Tennenholtz (2017) developed a best response linear regression predictor where two players compete and each gets a payoff depending on the proportion of the points he/she predicts more accurately than the other player. Dong et al. (2018) focused on the online version of the strategic classification algorithm. Chen et al. (2020b) developed a strategic-aware linear classifier to minimize the Stacelberg regret. Braverman and Garg (2020) modified the classifier to random ones. Moreover, Jagadeesan et al. (2021) added noise to standard strategic classification and modified the standard microfoundations into alternative microfoundations to let a portion of individuals be irrational and not have perfect knowledge about the decision maker’s policy. Tang et al. (2021) considered the setting where the decision maker only knew a subset of individuals’ actions. Levanon and Rosenfeld (2022) generalized strategic classification to situations where individuals and the decision maker have aligned interests. Perdomo et al. (2020); Izzo et al. (2021); Hardt et al. (2022); Jin et al. (2024) proposed and elaborated the concept of performative prediction where predictive decisions can influence the outcomes to predict. Izzo et al. (2021) proposed an algorithm performative gradient descent to compute performative optimal points. Hardt et al. (2022) defined performative power as a measure of how much a decision can change the population. This framework can formulate more general strategic classification settings. Liu et al. (2022) studied the situation where competitions between individuals are present in strategic classification. (Eilat et al., 2022) relaxed the assumption that individual best responses are independent of each other and proposed a robust learning framework based on a Graph Neural Network. Lechner and Urner (2022) proposed a novel loss function considering both the accuracy of the prediction rule and its vulnerability to strategic manipulation.
@@ A.2 Improvement with a label change
Another line of research takes improvement into account(Liu et al., 2018; Zhang et al., 2020b; Liu et al., 2020; Rosenfeld et al., 2020; Chen et al., 2020a; Haghtalab et al., 2020; Kleinberg and Raghavan, 2020; Alon et al., 2020; Miller et al., 2020; Shavit et al., 2020; Bechavod et al., 2021; Jin et al., 2022; Barsotti et al., 2022; Ahmadi et al., 2022a; Raab and Liu, 2021; Heidari et al., 2019). Liu et al. (2018, 2020); Zhang et al. (2020b); Rosenfeld et al. (2020); Ahmadi et al. (2022b) studied the conditions under which individuals will choose to improve their qualifications. Specifically, Liu et al. (2018) investigated how different decision rules (e.g. maxutil, fair) influence population qualification. Liu et al. (2020) modeled the improvement cost as a random variable and further pointed out that a subsidizing mechanism for individual costs can be beneficial for improving behaviors. Zhang et al. (2020b) studied the dynamic of population qualification under a partially observed Markov decision problem setting, where improvement probability is given as a parameter. Rosenfeld et al. (2020) proposed a Look-ahead regularization to directly penalize the drop of population qualification. Ahmadi et al. (2022b) proposed a common improvement capacity model and a individualized improvement capacity model to optimize social welfare and fairness while considering individual improvement.
There are other studies considering both strategic manipulation and improvement (Chen et al., 2020a; Haghtalab et al., 2020; Kleinberg and Raghavan, 2020; Alon et al., 2020; Miller et al., 2020; Shavit et al., 2020; Bechavod et al., 2021; Jin et al., 2022; Barsotti et al., 2022; Ahmadi et al., 2022a; Harris et al., 2022; Horowitz and Rosenfeld, 2023; Yan et al., 2023). Besides the works which have been mentioned in Sec. 1, (Barsotti et al., 2022) modeled strategic manipulation and improvement similarly with costs that differ within constant factors. The paper also did simulations where manipulation and improvement were present. (Ahmadi et al., 2022a) considered a general discrete model and a linear model where improvement and manipulation are both possible.
A.3 Machine learning fairness

### v2

@@ 4.2 Preferences as (dis)incentives for manipulation
Compared to the original strategic policy $\theta^{*}$, decision-makers can adjust preferences to disincentivize manipulation (i.e., $P_{M}(\theta)$ decreases) under certain scenarios. Specifically,
@@
    When either of the following is satisfied, and the decision-maker adjusts preferences by increasing $k_{1}$:
@@ 4.3 Preferences shape algorithmic fairness
The threshold shifts under adjusted preferences further allow us to compare these policies against a certain fairness measure. In this section, we consider strategic individuals from two social groups $\mathcal{G}_{a},\mathcal{G}_{b}$ distinguished by some protected attribute $S\in\{a,b\}$ (e.g., race, gender). Similar to Zhang et al. (2019; 2020b; 2022), we assume the protected attributes are observable and the decision-maker uses group-dependent threshold policy $\pi_{s}(x)=\textbf{1}(x\geq\theta_{s})$ to make decisions about $\mathcal{G}_{s},s\in\{a,b\}$. The optimal threshold for each group can be found by maximizing the utility associated with that group: $\max_{\theta_{s}}\mathbb{E}[R(D,Y)|S=s]$.
Fairness measure. We consider a class of group fairness notions that can be represented in the following form (Zhang et al., 2020a; Zhang & Liu, 2021):
@@
Thm. 4.6 identifies all scenarios under which a decision-maker can simultaneously promote fairness and disincentivize manipulation by simply adjusting $k_{1},k_{2}$. Otherwise, it is not guaranteed that both objectives can be achieved at the same time, as stated in Corollary 4.7.
A practical guideline for socially responsible decisions. The theoretical results in Thm. 4.5 give conditions under which the decision-maker can adjust preferences to disincentivize manipulation and encourage improvement, while the ones in Thm. 4.6 further sheds light on how the decision-maker can make explainable and socially responsible decisions under the unforeseeable strategic individual behavior: instead of adding separate regularizers to prevent manipulation or promote fairness, we show that the decision-maker may adjust their preferences in an interpretable way to disincentivize manipulation, incentivize improvement and promote fairness at the same time.
However, applying our theoretical results in practice needs access to several model parameters such as $q,\epsilon,C_{M},C_{I},F^{I}$ which can be estimated empirically. In App. B.5, we assume the decision-maker needs to estimate all parameters except the feature distribution for the qualified individuals $P_{X|Y}(x|1)$ and the qualification rate $\alpha$. Then we provide an estimation procedure for the parameters using controlled experiments on an experimental population [^4] Miller et al. (2020) already argued that designing policy for any strategic classification problem is a non-trivial causal modeling problem, thereby making the controlled experiments necessary for applications in practice.. With the model parameters, the decision-maker can derive accurate expressions for the decompositions $\phi_{1},\phi_{2},\phi_{3}$ and verify whether any condition in Thm. 4.5 and 4.6 holds. This enables the decision-maker to design a policy by adjusting preferences. For example, if estimated probabilities $q+\epsilon\geq 1$, then the decision-maker may train a policy with increased $k_{1}$ to disincentivize manipulation behavior (by Thm. 4.5, case 1).
@@ 5 Experiments
We conduct experiments on both synthetic Gaussian data and FICO score data (Hardt et al., 2016b)555code available at https://github.com/osu-srml/Unforeseeable-SC/tree/main.
FICO data (Hardt et al., 2016b).
FICO scores are widely used in the US to predict people’s credit worthiness. We use the preprocessed dataset containing the CDF of scores $F_{X|S}(x|s)$, qualification likelihoods $P_{Y|XS}(1|x,s)$, and qualification rates $\alpha_{s}$ for four racial groups (Caucasian, African American, Hispanic, Asian). All scores are normalized to $[0,1]$. Similar to Zhang et al. (2022), we use these to estimate the conditional feature distributions $P_{X|YS}(x|y,s)$ using beta distribution ${Beta}(a_{ys},b_{ys})$. The results are shown in Fig. 9. We assume the improved feature distribution $P^{I}(x)\sim{Beta}\big{(}\frac{a_{1s}+a_{0s}}{2},\frac{b_{1s}+b_{0s}}{2}\big{)}$ and $C_{M}-C_{I}\sim\mathcal{N}(0,0.25)$ for all groups, under which Assumption 2.2 and 2.1 are satisfied (see Fig. 8). We also considered other feature/cost distributions and observed similar results. Note that for each group $s$, the decision-maker finds its own optimal threshold ($\theta^{*}_{s}$ or $\theta^{*}_{s}(k_{i})$ or $\widehat{\theta}^{*}_{s}$) by maximizing the utility associated with that group, i.e., $\max_{\theta_{s}}\mathbb{E}[R(D,Y)|S=s]$.
@@
We also examine the individual best responses. Fig. 3 shows the manipulation probability $P_{M}(\theta)$ as a function of threshold $\theta$ for Caucasians and Asians (blue) versus African Americans and Hispanic (orange) when $q=0.3,\epsilon=0.5$. For all groups, there exists a unique $\theta_{max}$ that maximizes the manipulation probability. These are consistent with Thm. 2.3. We also indicate the manipulation probabilities under original strategic optimal thresholds $\theta_{s}^{*}$; The results indicate that original strategic optimal thresholds may cause the disadvantaged groups to have higher manipulation probabilities.
Table 2: Comparison between three types of optimal thresholds (FICO data). For Utility and $P_{M}$ values, the left value in parenthesis is associated with the advantaged group (Caucasian, Asian), while the right is for the disadvantaged group (African American, Hispanic). The first tabular is for Caucasian/African American while the second is for Asian/Hispanic.
@@
Threshold category Utility $P_{M}$ Unfairness (EqOpt)
Non-strategic $(0.726,0.427)$ $(0.115,0.322)$ 0.089
Original strategic $(0.734,0.448)$ $(0.055,0.161)$ 0.047
Adjusted strategic $(0.726,0.434)$ $(0.023,0.070)$ 0.022
Figure 3: $P_{M}(\theta)$ of Caucasian and African American (left plot) and of Asian and Hispanic (right plot).
Note that the scenario considered in Fig. 3 satisfies the condition 1.(ii) in Thm. 4.5, because the original strategic $\theta_{s}^{*}<\theta_{max}$ for both groups. We further conduct experiments in this setting to evaluate the impacts of adjusted preferences. We first adopt EqOpt as the fairness metric, under which $\mathbb{E}_{X\sim P_{s}^{\mathcal{C}}}[\textbf{1}(X\geq\widehat{\theta})]=F_{X|YS}(\theta|1,s)$ and the unfairness measure of group $\mathcal{G}_{a},\mathcal{G}_{b}$ can be reduced to $\big{|}F_{X|YS}(\theta|1,a)-F_{X|YS}(\theta|1,b)\big{|}$. Experiments for other fairness metrics are in App. C.1. The results are shown in Fig. 4, where dashed red and dashed blue curves are manipulation probabilities under non-strategic $\widehat{\theta}^{*}$ and strategic $\theta^{*}(k_{1})$, respectively. Solid red and solid blue curves are the actual utilities $U(\widehat{\theta}^{*})$ and $U(\theta^{*}(k_{1}))$ received by the decision-maker. The difference between the two dotted green curves measures the unfairness between Caucasians/African Americans or Asians/Hispanics. All weights are normalized such that $k_{1}=1$ corresponds to the original strategic policy, and $k_{1}>1$ indicates the policies with adjusted preferences. Results show that when condition 1(ii) in Thm. 4.5 is satisfied, increasing $k_{1}$ can simultaneously disincentivize manipulation ($P_{M}$ decreases with $k_{1}$) and improve fairness. These validate Thm. 4.5 and 4.6.
Figure 4: Impact of adjusted preferences: Caucasian and African American (left plot), Asian and Hispanic (right plot)
Figure 5: Impact of adjusted preferences (FICO data) when there is a Gaussian noise on $q$. The noises have 00 mean, and $0.05,0.1$ standard deviation from the left two plots to the right two plots.
@@
Tab. 2 compares the non-strategic $\widehat{\theta}^{*}$, original strategic ${\theta}^{*}$, and adjusted strategic ${\theta}^{*}(k_{1})$ when $k_{1,c}=k_{1,aa}=1.5$. It shows that decision-makers by adjusting preferences can significantly mitigate unfairness and disincentivize manipulation, with only slight decreases in utilities.
Robustness of results when $q,\epsilon$ are noisy.
We also present experiments to relax the assumption that the decision-maker knows $q,\epsilon$ exactly on the Caucasian/African groups. Instead, they only know $q+\delta$ or $\epsilon+\delta$ where $\delta$ is a Gaussian noise. We do $10$ rounds of simulations and produce plots with expectation and error bars similar to Fig. 4 (Fig. 6 shows the results with noisy $q$, while Fig. 6 shows the results with noisy $\epsilon$). The results show adjusting $k$ still works under noisy $q$ and $\epsilon$.
Gaussian Data.
@@
We also validate our theorems on synthetic data with Gaussian distributed $P_{X|YS}$ in App. C.2. Specifically, we examined the impacts of adjusting preferences on decision policies, individual’s best response, and algorithmic fairness. As shown in Fig. 19, 20 and Tab. 577, these results are consistent with theorems, i.e., adjusting preferences can effectively disincentivize manipulation and improve fairness. Notably, we considered all three scenarios in Thm. 4.5 when condition 1.(i) or 1.(ii) or 2 is satisfied. For each scenario, we illustrate the individual’s best response $P_{M}$ in Fig. 19 and show that manipulation can be disincentivized by adjusting preferences, i.e., increasing $k_{1}$ under condition 1.(i) or 1.(ii), or increasing $k_{2}$ under condition 2.
6 Conclusions & Limitations
@@
This paper proposes a novel probabilistic framework and formulates a Stackelberg game to tackle imitative strategic behavior with unforeseeable outcomes. Moreover, the paper provides an interpretable decomposition for the decision-maker to incentivize improvement and promote fairness simultaneously. The theoretical results depend on some (mild) assumptions and are subject to change when $\epsilon,q,C_{M},C_{I}$ change. Although we provide a practical estimation procedure to estimate the model parameters, it still remains a challenge to estimate model parameters accurately due to the expensive nature of doing controlled experiments. This may bring uncertainties in applying our framework accurately in real applications.
Broader Impact Statement
@@
Acknowledgement
This material is based upon work supported by the U.S. National Science Foundation under award IIS-2202699 and IIS-2416895, by OSU President’s Research Excellence Accelerator Grant, and grants from the Ohio State University’s Translational Data Analytics Institute and College of Engineering Strategic Research Initiative.
@@ A.1 Strategic classification
Generally, strategic behaviors can cause feature and label distribution of individuals to shift, which have long been closely related to concept drift (Lu et al., 2018), preference shift (Carroll et al., 2022), and algorithm recourse (Karimi et al., 2022). Strategic classification has been extensively studied since (Hardt et al., 2016a) formally modeled the interaction between individuals and a decision maker as a Stackelberg Game, and proposed a framework for strategic classification. While taking the individuals’ best response into account, the decision maker can make the optimal decision by anticipating strategic manipulation. During recent years, more complex models on strategic classification have been proposed (Ben-Porat & Tennenholtz, 2017; Dong et al., 2018; Braverman & Garg, 2020; Jagadeesan et al., 2021; Izzo et al., 2021; Ahmadi et al., 2021; Tang et al., 2021; Zhang et al., 2020b; 2022; Eilat et al., 2022; Liu et al., 2022; Lechner & Urner, 2022; Chen et al., 2020b; Xie & Zhang, 2024a). Ben-Porat & Tennenholtz (2017) developed a best response linear regression predictor where two players compete and each gets a payoff depending on the proportion of the points he/she predicts more accurately than the other player. Dong et al. (2018) focused on the online version of the strategic classification algorithm. Chen et al. (2020b) developed a strategic-aware linear classifier to minimize the Stacelberg regret. Tang et al. (2021) considered the setting where the decision maker only knew a subset of individuals’ actions. Levanon & Rosenfeld (2022) generalized strategic classification to situations where individuals and the decision maker have aligned interests. Lechner & Urner (2022) proposed a novel loss function considering both the accuracy of the prediction rule and its vulnerability to strategic manipulation. Eilat et al. (2022) relaxed the assumption that individual best responses are independent of each other and proposed a robust learning framework based on a Graph Neural Network. Horowitz et al. (2024) considered the self-selection problem in strategic classification where agents decide whether to participate in the system. Xie & Zhang (2024b) proposed a welfare-aware optimization framework in strategic learning settings.
Regarding randomness, Braverman & Garg (2020) proposed that randomized linear classifiers can be more robust to strategic behaviors. Moreover, Jagadeesan et al. (2021) added noise to standard strategic classification and modified the standard microfoundations into alternative microfoundations to let a portion of individuals be irrational and not have perfect knowledge about the decision maker’s policy. However, they just consider directly adding noise to the best response without modeling the unforeseeable and imitative nature of human agents’ behavior.
@@ A.2 Improvement with a label change
Another line of research takes improvement into account(Liu et al., 2018; Zhang et al., 2020b; Liu et al., 2020; Rosenfeld et al., 2020; Chen et al., 2020a; Haghtalab et al., 2020; Kleinberg & Raghavan, 2020; Alon et al., 2020; Miller et al., 2020; Shavit et al., 2020; Bechavod et al., 2021; Jin et al., 2022; Barsotti et al., 2022; Ahmadi et al., 2022a; Raab & Liu, 2021; Heidari et al., 2019; Somerstep et al., 2024). Liu et al. (2018; 2020); Zhang et al. (2020b); Rosenfeld et al. (2020); Ahmadi et al. (2022b) studied the conditions under which individuals will choose to improve their qualifications. Specifically, Liu et al. (2018) investigated how different decision rules (e.g. maxutil, fair) influence population qualification. Liu et al. (2020) modeled the improvement cost as a random variable and further pointed out that a subsidizing mechanism for individual costs can be beneficial for improving behaviors. Zhang et al. (2020b) studied the dynamic of population qualification under a partially observed Markov decision problem setting, where improvement probability is given as a parameter. Rosenfeld et al. (2020) proposed a Look-ahead regularization to directly penalize the drop of population qualification. Ahmadi et al. (2022b) proposed a common improvement capacity model and a individualized improvement capacity model to optimize social welfare and fairness while considering individual improvement. Jin et al. (2022); Chen et al. (2023) focused on designing subsidy mechanisms to incentivize improvement.
A.3 Studies considering both behaviors
There are other studies considering both strategic manipulation and improvement, but most of them modeled manipulation and improvement in a similar way(Chen et al., 2020a; Haghtalab et al., 2020; Kleinberg & Raghavan, 2020; Alon et al., 2020; Miller et al., 2020; Shavit et al., 2020; Bechavod et al., 2021; Jin et al., 2022; Barsotti et al., 2022; Ahmadi et al., 2022a; Harris et al., 2022b; Horowitz & Rosenfeld, 2023; Yan et al., 2023; Chen et al., 2023). Specifically, as illustrated in the abstract, the above-listed works all assume the agents can foresee the outcomes of their actions and they can change their features arbitrarily as long as the cost function permits. Moreover, these works modeled improvement and manipulation by assigning causal/non-causal features, which may not cover the practical cases in our paper such as college admission and job application. We further compare the differences between our model and causal strategic learning in App. B.1. Additionally, Perdomo et al. (2020); Izzo et al. (2021); Hardt et al. (2022); Jin et al. (2024) proposed and elaborated the concept of performative prediction where predictive decisions can influence the outcomes to predict. In general, this framework can model both manipulation and improvement but lacks interpretability.
There are only a few works considering both actions while incorporating randomness (Harris et al., 2022a; Bracale et al., 2024). Harris et al. (2022a) mainly focused on the situation where the decision-maker can choose from disclosing partial information of the model (namely, persuasion) and the agents estimate the best response using the partial information. Since the randomness completely comes from the information revealed by the decision-maker, this setting is different from ours where the randomness comes from the unforeseeable and imitative nature of agent improvement and manipulation. As a concurrent work, Bracale et al. (2024) considered a general setting where different actions incur random costs and lead to different feature-label distributions. However, they only focused on estimating the distribution map generally, which is tangential to our focus on disincentivizing manipulation, incentivizing improvement, and promoting fairness. Their contribution to estimating the distribution map may further facilitate the practical applications of our model.
A.4 Machine learning fairness